OpenStudy (anonymous):

Given f(x)=2x+3 and g(x)=1/2(x-3), find the each composition. I am confuse how to set then up.

OpenStudy (anonymous):

a composition is f(g(x)) nd g(f(x)), where you substitute the whole other function for x

OpenStudy (anonymous):

yeah,

OpenStudy (anonymous):

you start from the inside to the outside, but i still confuse how to set them up.

OpenStudy (amistre64):

f(g(x)) = 2(g(x)) + 3

OpenStudy (anonymous):

f(x)=2x+3 and g(x)=1/2(x-3), f(g(x) = 2(g(x)) + 3 f(g(x) = 2(1/2(x-3)) +3

OpenStudy (amistre64):

g(f(x)) = 1/2(f(x)-3)

OpenStudy (anonymous):

lol, amistre you fail

OpenStudy (amistre64):

:) how so?

OpenStudy (anonymous):

i just confuse, i need steps

OpenStudy (anonymous):

you just write something them leave.

OpenStudy (amistre64):

is there a time limt to wait around for a response? or am i allowed to come back after a response has been given? ;)

OpenStudy (amistre64):

i figure with 352 medals, i might be doing something right lol

OpenStudy (anonymous):

should we set one?

OpenStudy (anonymous):

WB, which steps do you get? That'd be a good start

OpenStudy (anonymous):

ok, so plz do step this time.

OpenStudy (anonymous):

the beginning,

OpenStudy (amistre64):

those are the steps; plug the one into the other; thats all there is to it.

OpenStudy (anonymous):

what you plug?

OpenStudy (amistre64):

f(x) = 2x+3 means that 'whatever' x equals, plug it in. f(g(x)) simply means that x = g(x)

OpenStudy (anonymous):

f(x)=2x+3 and g(x)=1/2(x-3), f(g(x) = 2(g(x)) + 3 f(g(x) = 2(1/2(x-3)) +3 g(f(x)) = 1/2(f(x)-3) g(f(x)) = 1/2(2x+3)-3) then simplify by working out how to remove the parentheses

OpenStudy (amistre64):

f(g(x)) = 2(g(x)) +3

OpenStudy (anonymous):

ah, got it. thx tonks

OpenStudy (anonymous):

all set, then I'm going to go now...jk ;)

OpenStudy (amistre64):

f(w) = 2(x) + 3 f(4) = 2(4) + 3 f(applesause) = 2(applesauce) + 3 f(g(x)) = 2(g(x)) + 3 lol

OpenStudy (amistre64):

...typed that lol

OpenStudy (anonymous):

wait, i have a similar problem, bvut much harder.

OpenStudy (anonymous):

the was just the start. and i got he answer, my prof did it on the board.

OpenStudy (amistre64):

Tonks has it ;)

OpenStudy (anonymous):

thx amistre

OpenStudy (anonymous):

WB, post it as a new question, it is easier to find that way

OpenStudy (anonymous):

f(x)=-7/2x -3 g(x)=-2x+6/7

OpenStudy (anonymous):

this one is much harder.

OpenStudy (anonymous):

WB, there's a new thread with this posted, check it out...